# A-OctoMap: An Adaptive OctoMap for Online Motion Planning

### v1

@@ Abstract
Traditional robotic motion planning methods often struggle with fixed resolutions in dynamically changing environments. To address these challenges, we introduce the A-OctoMap, an adaptive Octo-Tree structure that enhances spatial representation and facilitates real-time, efficient motion planning. This novel framework allows for dynamic space partitioning and multi-resolution queries, significantly improving computational efficiency and precision. Key innovations include a tree-based data structure for enhanced geometric processing, real-time map updating for accurate trajectory planning, and efficient collision detection. Our extensive testing shows superior navigation safety and efficiency in complex settings compared to conventional methods. A-OctoMap sets a new standard for adaptive spatial mapping in autonomous systems, promising significant advancements in navigating unpredictable environments.
@@ I-A Motivation
In the realm of robotics, particularly in motion planning, identifying safe zones in dynamic environments and effectively navigating through complex, narrow settings via optimized control techniques is a critical research area. The ability to swiftly and accurately define safe regions in such dynamic conditions is pivotal for enhancing trajectory planning efficiency and achieving optimal control.
In complex and cluttered environments filled with narrow passages, it becomes especially crucial to implement trajectory planning swiftly while maintaining effective tracking and real-time control over these trajectories (see [1, 2, 3]).
In the domain of spatial representation, occupancy grid mapping has been a foundational approach due to its straightforward representation of space[4]. However, this method is often limited by its fixed resolution, which constrains adaptability and necessitates a trade-off between detail and storage efficiency.
To address these challenges, we propose a new framework incorporating a novel tree-based data structure for dynamic space partition. This enhancement improves the identification of feasible solutions in congested environments. Our algorithm retains the capability for multi-resolution queries and augments data integration. Utilizing a tree-based data structure, it facilitates multi-threaded geometric processing of localized spatial segments post-partitioning, achieving effective storage and precise representation of spatial data with minimal computational expense. Moreover, our approach ensures the real-time updating of the map in dynamic environments while preserving critical geometric information, which is vital for dynamic path planning and collision detection.
@@ I-B1 Volumetric Representation of Space
Traditional volumetric reconstruction approaches in robotics have primarily leveraged voxel-based methods, such as voxel-based techniques and hierarchical structures like Octo-Trees[5]. These methods have been favored for their rapid execution, which is crucial for on-the-fly applications in dynamic environments. However, they often fall short when it comes to capturing fine surface details, as they usually sacrifice spatial resolution for computational efficiency.
Despite the speed advantages of voxel-based strategies for 3-D mapping, their lack of granularity often fails in high-fidelity tasks where distinguishing between traversable and obstructed spaces is crucial. Advanced techniques like the Truncated Signed Distance Field (TSDF) and the Euclidean Signed Distance Field (ESDF) offer better distance metrics for motion planning but struggle with data coarseness, limiting resolution and precision in collision avoidance[6, 7]. Additionally, TSDF methods are memory-intensive and computationally challenging to update in real-time for larger environments[8]. In contrast, the emergence of Neural Radiance Fields (NeRF) through deep learning innovations has significantly enhanced autonomous 3-D scene reconstruction, excelling in capturing detailed object surfaces and facilitating sophisticated view planning essential for Simultaneous Localization and Mapping (SLAM)[9, 10]. However, while promising, the computational demands and potential latency in real-time rendering by these neural representation-based methods pose challenges, particularly in dynamic environments requiring high-speed maneuvers.
@@ I-B2 Point Cloud Downsampling
The development of point cloud downsampling techniques has primarily focused on balancing efficiency and the preservation of geometric features. Traditional strategies like Farthest Point Sampling (FPS) and Random Sampling (RS) each have their strengths and weaknesses. RS, which involves the random selection of points, is straightforward and fast but fails to preserve geometric features and ensure uniform coverage, resulting in significant information loss [11, 12]. In contrast, FPS iteratively selects the farthest points from those already chosen, ensuring better geometric feature preservation compared to random sampling. However, its computational intensity restricts its use in real-time applications [13].
Voxel-based sampling (VBS) divides space into a grid of voxels and retains a single representative point per voxel, either the gravity centroid or the original point closest to the centroid. This method offers high computational efficiency suitable for real-time applications and can be combined with RS and FPS strategies, including Uniform Voxel Sampling (UVS) and Voxelized Farthest Points Sampling (VFPS)[14, 15]. However, its fixed voxel size limits adaptability, leading to detail loss in high-density regions, and the use of centroids to represent all points within a voxel results in constant local density.
3-D Edge-Preserving Sampling (3DEPS) has emerged to address these limitations in detail loss[15]. Inspired by sketching techniques, it captures the essence of complex 3-D shapes by emphasizing their sharp edges, providing a more effective representation of intricate objects. 3DEPS utilizes the 3-D Surface Boundary Filter (SBF) to separate the point cloud into edge points and internal points[16]. By adjusting the ratio of these two types of points, a new point cloud is reconstructed, ensuring a balance between preserving edge details and overall point distribution. However, the ratio of edge points to the total number of points is a parameter that must be experimentally tuned, adding complexity to the process.
@@ I-B3 Path Planning Algorithms
In motion planning, various methods can be employed to ensure effective collision-free paths. Sampling-based algorithms, such as Rapidly-exploring Random Trees (RRT) and Probabilistic Roadmaps (PRM), achieve impressive results in high-dimensional, complex spaces by randomly expanding searches within the feasible space to find a viable path[17, 18]. The efficiency of these methods heavily relies on the techniques used for sampling points in free regions and the approach to nearest neighbor searches.
Grid search-based algorithms (like Dijkstra, A*, and their variants) offer resolution completeness ensuring the determination of the shortest path between nodes in grid-based searches. A*, an extension of the Dijkstra algorithm, utilizes a heuristic approach to estimate the overall path cost, thereby outperforming the Dijkstra algorithm. Jump Point Search (JPS) [19], the optimized algorithm of A*, was proposed for uniform grids to accelerate search speeds by eliminating certain nodes in the grid based on the setup of jump points. However, all these method are heavily limited by its fixed resolution, which constrains adaptability and necessitates a trade-off between detail and storage efficiency.
@@ I-C Contributions
    Dynamic space representation: Our dynamic Octo-Tree achieves high-precision segmentation of free spaces and obstacles, effectively addressing the complexities of dynamic environments.
@@
    Spatial representation fidelity: This method supports refined spatial operations and downsampling.
@@
    Efficient collision detection: By parallelizing maximal map polyhedronization and surface triangulation, we achieve efficient collision detection in dynamic environments.
  4. 4.
    Integration with navigation methods: This framework accelerates the identification of feasible solutions for grid-based navigation methods (e.g., JPS), thereby enhancing overall navigation efficiency.
@@ II Preliminaries
In this section, we introduce some definitions and results on Adaptive OctoMap and Jump Point Search (JPS).
II-A Adaptive OctoMap Node
The attribution of a tree node in the Adaptive OctoMap is present as Data Structure 1
Algorithm 1 Data Structure 1: OctreeNode
@@ II-A Adaptive OctoMap Node
A tree $T$ can be represented as $T=(V,E)$ by a set of nodes $V$ and a set of edges $E$:
$V=\{v_{1},v_{2},\dots,v_{n}\},\quad E=\{(v_{i},v_{j})\mid v_{i},v_{j}\in V\}$ (1)
Each node $v_{i}$ is defined as:
$v_{i}=(\text{value},\text{parent},\text{children})$ (2)
where value represents the identifier of Point Cloud, parent represents the parent node ($\emptyset$ if it is the root), and children represents the set of child nodes.
In this paper, we consider the point cloud as the map storage resource, represented as follows:
$\mathcal{P}=\{\mathbf{p}_{i}\mid\mathbf{p}_{i}\in\mathbb{R}^{d},\,i=1,2,\ldots,N,\,d\in\{2,3\}\}$ (3)
The point cloud $\mathcal{P}$ is represented as a set containing $N$ points, where each point $\mathbf{p}_{i}$ can be either two-dimensional or three-dimensional.
The attributes SplitBoundary and NodeBoundary define the boundaries of sub-nodes in an Octree:
  1. 1.
    SplitBoundary: refers to the boundaries of each sub-node after a node is divided along its axes.
  2. 2.
    NodeBoundary: represents the actual maximum and minimum values along each axis for the sub-nodes after complete bisection.
II-B Adaptive OctoMap Tree
The attribution of a tree in the Adaptive OctoMap is present as Data Structure 2:
Algorithm 2 Data Structure 2: Octree
1:  Struct OcTree
@@ II-B Adaptive OctoMap Tree
In this paper, we introduce the concept of Minimum Controllable Region (MCR), which describes the smallest area within which a robot can perform precise control given its control accuracy. This value quantifies the minimal space in which the robot can operate effectively, considering its control precision.
Definition 2.
The MCR is defined by
where $\mathbf{p}=(x_{1},x_{2},\ldots,x_{d})$ is a point in space, $\mathbf{p}_{0}=(x_{0,1},x_{0,2},\ldots,x_{0,d})$ is the reference point (typically the robot’s current position), $\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}=\max\{|x_{1}-x_{0,1}|,|x_{2}-x_{0,2}|,\ldots,|x_{d}-x_{0,d}|\}$ is the $infinity$ norm (i.e., the maximum absolute difference in any dimension), and the $\epsilon_{\max}$ is the estimated upper bound of error.
To determine the depth of a tree, given the exploration range $L$ along a particular axis (which is determined by the upper bound of the maximum system error $\epsilon_{\max}$ along the same axis, i.e., the range can be greater than or equal to $\epsilon_{\max}$), we use the following formula:
[…]
II-C Uniform Grid Map
The attribution of a uniform grid map is present as Data Structure 3
Algorithm 3 Data Structure 3: uniform grid map
1:  Struct UniformGridMap
2:   array <int, d>grid_map_amount
3:   multi_array <bool, d>map
Theorem 1 (Pigeonhole principle).
For natural numbers $k$ and $m$, if $n=km+1$ objects are distributed among $m$ sets, the pigeonhole principle asserts that at least one of the sets will contain at least $k+1$ objects[20].
We conclude that if a grid cell in a uniform grid map has a side length of $n_{s}$, and there exists a gap of at least $n_{s}$ meters within this map, then the boundary of at least one cell must pass through this gap.
II-D Jump Point Search
A graph search (path planning) algorithm can then be used to identify a valid path that avoids collisions within the grids occupied by obstacles. JPS is suitable for rapid path planning. we define a path, denoted as $\pi=\left\langle n_{i},\ldots,n_{j}\right\rangle$, as a sequence of nodes or states extending from node $i$ to node $j$. The length (or cost) of this path is represented by $\text{len}(\pi)$. We introduce the notation $x$ to refer to a specific node, $p(x)$ to denote the parent node of $x$, and $\text{neighbours}(x)$ to describe the set of nodes adjacent to $x$. Additionally, the node $n$ is defined as an element within the set of $\text{neighbours}(x)$. The expression $\pi\backslash x$ signifies that node $x$ is omitted from the path.
In a 2-D map, each node may have up to 8 neighbors. The cost incurred when moving straightly (horizontally or vertically) to a traversable neighbor (not occupied by obstacles) is 1, while moving diagonally has a cost of $\sqrt{2}$. For straight moves, if node $n$ can be reached from $x$’s parent $p(x)$ with the less or equal cost without passing through $x$, i.e., $len(\left\langle p(x),\dots,n\right\rangle\backslash x)\leq len(\left\langle p(x),x,n\right\rangle)$, JPS will prune this neighbor $n$ from exploration. In diagonal moves, the path excluding $x$ must be strictly shorter, $len(\left\langle p(x),\dots,n\right\rangle\backslash x)<len(\left\langle p(x),x,n\right\rangle)$, leading to the pruning of node $n$. The natural neighbors of $x$ are those remaining after pruning, assuming $\text{neighbours}(x)$ does not include obstacles. Force Neighbors refer to specific neighboring nodes that may not lie directly on the current path but must be examined to ensure the identification of the shortest path, which must satisfy with two condtions. 1) n is not a natural neighbour 2) $len(\left\langle p(x),x,n\right\rangle)<len(\left\langle p(x),\dots,n\right\rangle\backslash x)$.A node qualifies as a jump point if it is either the initial or terminal node, has a forced neighbor, or, in the case of diagonal movement, leads directly to another jump point. as detailed in [19] and [1].
Starting from initial node, the algorithm recursively searches for new jump points in every direction until the terminal node is found or a dead end is reached in that direction. Upon reaching the terminal node, it backtracks through the jump points to construct the shortest path, $\pi=\left\langle n_{0},n_{1},\dots,n_{E}\right\rangle$, from start to finish. Path searching with jump point pruning has been proven to be cost-optimal in [19].
III Methodology
III-A Tree build
We derive the tree depth from Equation (5), ensuring the tree fully covers the entire map. As described in Algorithm 1 in III-A, the method relies on a multi-level tree structure. This process involves partitioning each level through bisection along different dimensions, ensuring precise allocation of points to their corresponding child nodes. Additionally, each node’s boundary is defined by selecting the extremal values (maximum and minimum) in each dimension after the bisection process. These values, along with the node’s boundary, are stored separately in SplitBoundary and NodeBoundary. The free region within each node can be indirectly represented by these two boundaries at each level. This recursive process is repeated until a leaf node is reached, at which point the current node’s address is pushed into the vector $\mathcal{L}$.
Algorithm 4 Algorithm 1: Octree Build

### v2

# A-OctoMap: An Adaptive OctoMap for Online path planning

@@ Abstract
Downsampling and path planning are essential in robotics and autonomous systems, as they enhance computational efficiency and enable effective navigation in complex environments. However, current downsampling methods often fail to preserve crucial geometric information while maintaining computational efficiency, leading to challenges such as information loss during map reconstruction and the need to balance precision with computational demands. Similarly, current graph-based search algorithms for path planning struggle with fixed resolutions in complex environments, resulting in inaccurate obstacle detection and suboptimal or failed pathfinding. To address these issues, we introduce an adaptive OctoMap that utilizes a hierarchical data structure. This innovative approach preserves key geometric information during downsampling and offers a more flexible representation for pathfinding within fixed-resolution maps, all while maintaining high computational efficiency. Simulations validate our method, showing significant improvements in reducing information loss, enhancing precision, and boosting the computational efficiency of map reconstruction compared to state-of-the-art methods. For path planning, our approach enhances Jump Point Search (JPS) by increasing the success rate of pathfinding and reducing path lengths, enabling more reliable navigation in complex scenes.
@@ I-A Motivation
Downsampling and path planning are essential in robotics and autonomous systems, improving computational efficiency and enabling reliable navigation in complex environments. Downsampling simplifies sensor data, allowing real-time processing while preserving key map features such as obstacles. Path planning ensures safe and optimal routes by avoiding collisions. Many downsampling approaches use hierarchical structures like voxel grids [1] and OctoTrees to achieve efficiency. OctoMap [2], a probabilistic mapping framework built on OctoTree, incorporates occupancy information for 3-D mapping. However, such methods often struggle to capture fine geometric details. In contrast, geometry-preserving techniques offer better detail retention but suffer from lower computational efficiency [3, 4].
In the domain of path planning, graph-based search algorithms [5, 6] have long served as a foundational approach due to their straightforward representation of space. However, these methods are constrained by their fixed resolution, limiting their adaptability to environmental complexities. A fixed resolution often fails to capture accurate obstacle boundary details or narrow passages, making it difficult to find globally optimal paths and potentially leading to path finding failures.
To address the limitations of hierarchical structures in downsampling and fixed-resolution mapping for path planning, we utilize an optimized OctoMap, a hierarchical tree-based data structure that represents point cloud clusters and obstacles at multiple scales. This approach allows downsampling to preserve essential geometric features critical for accurate map reconstruction and provides a more flexible representation of obstacles, enhancing the optimality and feasibility of pathfinding compared to traditional fixed-resolution maps.
@@ I-B1 Volumetric Representation of Space
Traditional volumetric reconstruction approaches in robotics have primarily leveraged voxel-based methods, such as voxel-based techniques and hierarchical structures like OctoTrees[2]. These methods have been favored for their rapid execution, which is crucial for on-the-fly applications in dynamic environments. However, they often fall short when it comes to capturing fine surface details, as they usually sacrifice spatial resolution for computational efficiency.
While voxel-based 3D mapping is computationally efficient, it often lacks the granularity required for high-fidelity tasks such as precise obstacle avoidance. Methods like Truncated Signed Distance Fields (TSDF) and Euclidean Signed Distance Fields (ESDF) improve distance metrics [3, 4], but still suffer from data coarseness and high memory costs, limiting their scalability [7]. Neural Radiance Fields (NeRF) offer detailed 3D reconstructions [8, 9], yet their computational demands hinder real-time use in dynamic settings. In contrast, our method adaptively refines map resolution only in geometrically complex regions, preserving fine object boundaries without uniformly increasing memory or computation, unlike traditional OctoMap or fixed-resolution voxels.
@@ I-B2 Point Cloud Downsampling
Point cloud downsampling techniques aim to balance efficiency with the preservation of geometric features. Traditional methods like Random Sampling (RS) and Farthest Point Sampling (FPS) each have distinct trade-offs. RS is simple and fast, selecting points randomly, but often fails to preserve geometric structure or ensure uniform coverage, leading to significant information loss [10, 11]. In contrast, FPS iteratively selects the farthest points from the chosen set, offering better geometric fidelity, but its high computational cost limits real-time applicability [12].
Voxel-based sampling (VBS) divides space into a grid of voxels and retains a representative point per voxel, either the centroid or the nearest original point [1, 13]. This method offers high computational efficiency suitable for real-time applications and can be combined with RS and FPS strategies, such as Uniform Voxel Sampling (UVS) and Voxelized Farthest Point Sampling (VFPS). However, its fixed voxel size limits adaptability, leading to detail loss in high-density regions. Moreover, representing all points in a voxel with a single centroid results in constant local density and loss of geometric variation.
3-D Edge-Preserving Sampling (3DEPS) addresses these limitations by emphasizing sharp features [13]. Inspired by sketching techniques, it captures complex 3-D shapes more effectively by applying a 3-D Surface Boundary Filter (SBF) to separate edge points from internal points [14]. A new point cloud is then reconstructed by adjusting the edge-to-internal point ratio, balancing detail preservation and overall distribution. However, this ratio requires empirical tuning, increasing system complexity. We improve downsampling by constructing convex hulls from node-boundary extremal points and recursively eliminating interior points. This preserves sharp features and spatial structure more effectively than voxel or random sampling, while supporting real-time performance through parallel processing.
@@ I-B3 Path Planning Algorithms
In path planning, various methods can be employed to ensure effective collision-free paths. Sampling-based algorithms, such as Rapidly-exploring Random Trees (RRT) and Probabilistic Roadmaps (PRM), achieve impressive results in high-dimensional, complex spaces by randomly expanding searches within the feasible space to find a viable path[15, 16]. The efficiency of these methods heavily relies on the techniques used for sampling points in free regions and the approach to nearest neighbor searches.
Grid search-based algorithms, such as Dijkstra, A*, and their variants, offer resolution completeness, ensuring the shortest path can be found in grid-based searches. A*, an extension of Dijkstra, introduces heuristics to estimate path cost and improves search efficiency. Jump Point Search (JPS) [6], a grid-optimized variant of A*, accelerates search by pruning unnecessary nodes based on jump point logic. However, these methods are limited by fixed grid resolution, which restricts adaptability and forces a trade-off between detail and memory usage. Rather than modifying JPS itself, we enhance path planning by constructing an adaptive uniform grid aligned with OctoTree boundaries, capturing obstacle edges more precisely. This improves feasibility in narrow or irregular regions, where fixed-resolution grids may fail, leading to higher success rates and shorter paths. To address these challenges, we propose an adaptive grid mapping framework that integrates enhanced OctoMap structures with uniform grid projection and dynamic spatial refinement, enabling efficient environment representation while preserving obstacle boundaries and supporting real-time planning.
@@ I-C Contributions
    We introduce an adaptive OctoMap with multiscale segmentation, dynamically adjusting resolution based on geometric and control precision requirements. This allows fine-grained representation in complex regions without global overhead.
@@
    Our method enables parallel downsampling via local convex hull approximation, preserving sharp features while reducing point density and maintaining real-time feasibility.
@@
    Our adaptive grid mapping framework integrates with JPS, allowing it to operate on geometry-aware grids that improve both success rate and path length in complex environments.
@@ II Preliminaries
In this section, we provide background information on uniform-grid OctoMap and Jump Point Search (JPS).
II-A Uniform-grid OctoMap
Figure 1: OctoTree map with hierarchical voxels.
Figure 2: Uniform grid map based on Data Structure 1.
Although OctoMap is typically implemented as an adaptive OctoTree-based map, it can also be simplified into a uniform-resolution grid by fixing the voxel size and disabling recursive subdivision. This simplified version retains the OctoMap framework and data access methods, while allowing efficient indexing like a uniform grid. We refer to this representation as the Uniform Grid OctoMap, which is illustrated in Fig. 2 and formalized in Data Structure 1.
Algorithm 1 Data Structure 1: Uniform Grid OctoMap
1:  Struct UniformGridMap
2:   array <int, d>grid_map_amount
3:   multi_array <bool, d>map
Figure  2 illustrates a traditional OctoTree structure, which recursively subdivides 3D space into hierarchical voxels. While this representation enables adaptive resolution and memory efficiency, it involves complex tree traversal during planning. In contrast, figure 2 shows a uniform grid map, where the environment is discretized into a fixed-resolution grid. This corresponds to Data Structure 1, in which the map is encoded using a fixed-size integer array to define grid dimensions (array<int, d>) and a multi-dimensional boolean array (multi_array<bool, d>) to indicate occupancy. The uniform grid structure offers faster access and simpler indexing, making it more suitable for efficient grid-based planners such as JPS.
Theorem 1 (Pigeonhole principle [17]).
For natural numbers $k$ and $m$, if $n=km+1$ objects are distributed among $m$ sets, the pigeonhole principle asserts that at least one of the sets will contain at least $k+1$ objects.
We apply Thm. 1 to the problem of uniform grid coverage. In a uniform grid map with fixed cell size $n_{s}$, if a gap of width at least $n_{s}$ exists in the environment, then by the pigeonhole principle, at least one grid cell boundary must intersect this gap. This ensures that any obstacle-free passage of sufficient width will be captured by the grid structure, and thus, it can be considered during planning. Therefore, the use of uniform-resolution grids is sufficient to detect narrow but traversable gaps, as long as the gap size exceeds the grid resolution.
II-B Jump Point Search
A graph search algorithm can be used to find a valid path that avoids collisions with obstacle-occupied grids. Jump Point Search (JPS) is particularly suited for rapid path planning. We define a path $\pi=\langle n_{i},\ldots,n_{j}\rangle$ as a sequence of nodes from $n_{i}$ to $n_{j}$, with its length (or cost) denoted by $\text{len}(\pi)$. Let $x$ be a node, $p(x)$ its parent, and $\text{neighbours}(x)$ the set of adjacent nodes. A specific neighbor $n$ is an element of $\text{neighbours}(x)$, and $\pi\setminus x$ denotes the path $\pi$ with node $x$ omitted.
In a 2-D map, each node may have up to 8 neighbors. Moving straight (horizontally or vertically) to a traversable neighbor costs 1, while diagonal moves incur a cost of $\sqrt{2}$.
For straight moves, a neighbor $n$ is pruned if it can be reached from $x$’s parent $p(x)$ with less or equal cost without passing through $x$, i.e.,
$\text{len}(\langle p(x),\dots,n\rangle\setminus x)\leq\text{len}(\langle p(x),x,n\rangle).$
For diagonal moves, the pruning condition is stricter:
$\text{len}(\langle p(x),\dots,n\rangle\setminus x)<\text{len}(\langle p(x),x,n\rangle).$
The natural neighbors of $x$ are those remaining after pruning, assuming $\text{neighbours}(x)$ excludes obstacles. Forced neighbors refer to nodes that are not natural neighbors but must be explored to ensure optimality. A node $n$ is a forced neighbor if:
  1. 1.
    $n$ is not a natural neighbor; and
  2. 2.
    $\text{len}(\langle p(x),x,n\rangle)<\text{len}(\langle p(x),\dots,n\rangle\setminus x)$.
A node qualifies as a jump point if it is the start or goal node, has a forced neighbor, or (in diagonal movement) leads directly to another jump point, as defined in [6, 18]. Starting from initial node, the algorithm recursively searches for new jump points in every direction until the terminal node is found or a dead end is reached in that direction. Upon reaching the terminal node, it backtracks through the jump points to construct the shortest path, $\pi=\left\langle n_{0},n_{1},\dots,n_{E}\right\rangle$, from start to finish. Path searching with jump point pruning has been proven to be cost-optimal in [6]. JPS has also been shown to efficiently generate safe paths in environments with clustered dynamic obstacles, guiding the system effectively [19].
III OctoMap-Based Adaptive Grid Mapping Framework
This section presents the proposed adaptive grid mapping framework, which builds upon the standard OctoMap representation to enable geometry-aware refinement and planning-oriented grid generation. At its core, we introduce a hierarchical data structure, referred to as the Adaptive OctoMap Tree, which supports recursive space decomposition while preserving local obstacle boundary information.
III-A TreeNode: Node Structure for Adaptive OctoMap Tree
Each node in the tree represents a localized region of space and maintains geometric and topological attributes that support adaptive partitioning. In contrast to traditional OctoMap nodes, our node structure includes a local point cloud segment and two axis-aligned bounding boxes: one representing the node’s physical boundary and the other capturing the position of the most recent split. The complete definition of the tree node structure is provided in Data Structure 2.
Algorithm 2 Data Structure 2: Node representation in the Adaptive OctoMap Tree
@@ III-A TreeNode: Node Structure for Adaptive OctoMap Tree
Each internal node subdivides its region into child subregions based on bisection along coordinate axes. The bisection information is encoded in two key fields:
  • SplitBoundary: the location of splitting planes along each axis.
  • NodeBoundary: the spatial extent (min/max bounds) of the node region.
These attributes are used during tree construction and map conversion to precisely identify geometric boundaries of obstacles and free space. To determine whether a node should be further subdivided, we introduce the concept of the Minimum Controllable Region (MCR). This concept establishes a principled lower bound on spatial resolution based on the agent’s ability to traverse narrow regions. The MCR defines the minimal region width required for safe navigation, and serves as a stopping condition for recursive tree expansion.
Definition 1 (Minimum Controllable Region (MCR)).
The MCR centered at a reference point $\mathbf{p}_{0}\in\mathbb{R}^{d}$ is defined as
where $\epsilon_{\max}$ denotes the maximum allowable deviation in any dimension.
Here, $\|\cdot\|_{\infty}$ denotes the infinity norm, i.e., the maximum absolute difference across dimensions. The value $\epsilon_{\max}$ is typically determined by the system’s localization uncertainty, physical footprint, or clearance requirement. This region characterizes the smallest space where an agent can reliably operate without further subdivision. To determine the required tree depth for a given axis-aligned map size $L$, we use:
where $p$ is the MCR size along that axis, and $k>1$ is a safety scaling factor (e.g., $k=2$) to ensure sufficient resolution.
III-B Adaptive OctoMap Tree Structure
Building on the TreeNode definition, we now formalize the structure of the entire adaptive tree used for recursive spatial subdivision. To enable traversal, refinement, and global reasoning, the tree is represented both structurally as a hierarchical object and formally as a graph.
Formally, we define the tree as a directed acyclic graph (DAG) where each node stores geometric, hierarchical, and relational information. This abstraction supports efficient representation and manipulation of map regions during adaptive mapping and planning.
@@ III-B Adaptive OctoMap Tree Structure
An adaptive OctoMap Tree $T$ is defined as a graph $T=(V,E)$ where:
  • $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is the set of nodes,
  • $E=\{(v_{i},v_{j})\mid v_{j}\text{ is a child of }v_{i}\}$ represents parent-child edges.
[…]
To support efficient access and recursive processing, we define the tree using the structure shown in Data Structure 3. It includes a root node and a tree depth, as determined by the MCR criterion in Eqn. (2).
Algorithm 3 Data Structure 3: OctoTree
1:  Struct OctoTree
@@
III-C Tree Construction
With the structure of the adaptive OctoMap Tree formally defined, we now describe the process of constructing this tree from a raw point cloud input. The goal is to recursively subdivide space based on geometric content and the MCR criterion until a resolution limit is reached. The tree construction process is recursive. Starting from a root node, the point cloud is progressively divided into smaller subregions. For each node, we first check if it satisfies the termination condition based on depth or obstacle sparsity. If not, we bisect the region along its longest axis, update the SplitBoundary and NodeBoundary, allocate points to child nodes, and repeat the process. The recursion ends when the region size reaches the MCR threshold or the depth limit is met. The full procedure is detailed in Algorithm 1.
Algorithm 1 Algorithm 1: OctoTree Construction